welch construction has found that the value of the company tractor is decreasing at a rate give by v(t)= -390te-0.012t^2, where t represents the number of years after purchase. if the company purchased the tractor for $17650, what is the value of the tractor 3 years after purchase


1
Expert's answer
2021-06-06T15:42:59-0400
"U(t)=\int v(t)dt=\int-390te^{-0.012t^2}dt"

"=-16250e^{-0.012t^2}+C"

Given "U(0)=17650"


"17650=-16250+C=>C=33900"

"U(t)=-16250e^{-0.012t^2}+33900"


"U(3)=-16250e^{-0.012(3)^2}+33900"

"U(3)=\$19313.55"